OpenStudy (anonymous):

How do I use the properties of logarithms to rewrite 1/2log y p^3q^3-2/3log y p^4q^3

OpenStudy (anonymous):

use parentheses please

OpenStudy (anonymous):

\[(1/2) \log _{y} (p ^{3}q ^{4})-(2/3) \log _{y}(p ^{4}q ^{3})\]

OpenStudy (anonymous):

okay, use the rules x log a = log a^x and log a - log b = log(a/b)

OpenStudy (anonymous):

what about the base y ?

OpenStudy (anonymous):

shouldn't it be y^2/3 ?

OpenStudy (anonymous):

it is constant in both terms.

OpenStudy (anonymous):

no. you can manipulated logarithms with the same base in the ways I indicated.

OpenStudy (anonymous):

is that all what do ido with the p and q's